OpenStudy (chris215):

Find the range for f(x) = x2 + 1, for x < 0. y ≥ 1 y > 1 y < 1 y ≤ 1 A?

OpenStudy (freckles):

\[x<0 \\ \text{ multiply } x \text{ on both sides } x \cdot x >0 \cdot x \\ \text{ do you know why I change the direction of the inequality ? }\]

OpenStudy (freckles):

hint: what does x<0 mean in words

OpenStudy (chris215):

x less than 0?

OpenStudy (freckles):

or x is negative if you multiply both sides of an inequality by a negative number you have to change the direction of the inequality

OpenStudy (freckles):

\[x<0 \\ x \cdot x> 0 \cdot x \\ x^2 >0 \\ \text{ now add 1 on both sides } x^2+1>0+1\]

OpenStudy (freckles):

\[x^2+1>1 \]

OpenStudy (freckles):

remember x^2+1 is y since we had f(x)=x^2+1

OpenStudy (chris215):

ohh ok thank you!

OpenStudy (freckles):

np